OpenStudy (anonymous):

One number is seven more than twice another. If their sum is decreased by nine, the result is thirteen. Find the numbers.

OpenStudy (anonymous):

let the number be x,let the other number be y is means =

OpenStudy (anonymous):

please show me i don not understand..

OpenStudy (anonymous):

\[x + 7 = 2y \Rightarrow y = \frac{x+7}{2}\]\[x + y - 9 = 13 \Rightarrow x + y = 21 \Leftrightarrow x + \frac{x+7}{2} = 21\]

OpenStudy (anonymous):

solve for x, then plug in the value you got into the formula for y

OpenStudy (anonymous):

so what the length and whats the width